# Flow Matching Policy Gradients

### v1

@@ 1 Introduction
We theoretically analyze FPO’s correctness and empirically validate its performance across a diverse set of tasks. These include a GridWorld environment, 10 continuous control tasks from MuJoCo Playground Zakka et al. (2025), and high-dimensional humanoid control—all trained from scratch. FPO demonstrates robustness across tasks, enabling effective training of flow-based policies in high-dimensional domains. We probe flow policies learned in the toy GridWorld environment and find that on states with multiple possible optimal actions, it learns multimodal action distributions, unlike Gaussian polices. On humanoid control tasks, we show that the expressivity of flow matching enables single-stage training of under-conditioned control policies, where only root-level commands are provided. In contrast, standard Gaussian policies struggle to learn viable walking behaviors in such cases. This highlights the practical benefits of the more powerful distribution modeling enabled by FPO. Finally, we discuss limitations and future work.
@@ 3.4 Estimating the FPO Ratio with Flow Matching
$\hat{r}^{\text{FPO}}_{\theta}(\tau,\epsilon)=\exp(\ell_{\theta_{\text{old}}}(\tau,\epsilon)-\ell_{\theta}(\tau,\epsilon)),$ (17)
@@
$\nabla_{\theta}\hat{r}^{\text{FPO}}_{\theta}(\tau,\epsilon)=-\hat{r}^{\text{FPO}}_{\theta}(\tau,\epsilon)\nabla_{\theta}\ell_{\theta}(\tau,\epsilon).$ (19)
Since the gradient operator commutes with expectation, the gradient term above is unbiased:
$\mathbb{E}_{\tau,\epsilon}[-\nabla_{\theta}\ell_{\theta}(\tau,\epsilon)]=-\nabla_{\theta}\mathcal{L}_{\theta}^{w}{}(a_{t})=\nabla_{\theta}\text{ELBO}_{\theta}(a_{t}).$ (20)
In other words, gradient estimates are directionally unbiased even with worst-case overestimation of ratios. Our experiments are consistent with this result; we observe empirically in Section 4.2 that FPO can be trained to outperform Gaussian PPO even with $N_{\text{mc}}=1$ though additional samples help.

### v2

@@ 1 Introduction
We theoretically analyze FPO’s correctness and empirically validate its performance across a diverse set of tasks. These include a GridWorld environment, 10 continuous control tasks from MuJoCo Playground Zakka et al. (2025), and high-dimensional humanoid control—all trained from scratch. FPO demonstrates robustness across tasks, enabling effective training of flow-based policies in high-dimensional domains. We probe flow policies learned in the toy GridWorld environment and find that on states with multiple possible optimal actions, it learns multimodal action distributions, unlike Gaussian policies. On humanoid control tasks, we show that the expressivity of flow matching enables single-stage training of under-conditioned control policies, where only root-level commands are provided. In contrast, standard Gaussian policies struggle to learn viable walking behaviors in such cases. This highlights the practical benefits of the more powerful distribution modeling enabled by FPO. Finally, we discuss limitations and future work.
@@ 3.4 Estimating the FPO Ratio with Flow Matching
$\hat{r}^{\text{FPO}}_{\theta}(\tau,\epsilon)=\exp(\ell^{w}_{\theta_{\text{old}}}(\tau,\epsilon)-\ell^{w}_{\theta}(\tau,\epsilon)),$ (17)
@@
$\nabla_{\theta}\hat{r}^{\text{FPO}}_{\theta}(\tau,\epsilon)=-\hat{r}^{\text{FPO}}_{\theta}(\tau,\epsilon)\nabla_{\theta}\ell^{w}_{\theta}(\tau,\epsilon).$ (19)
Since the gradient operator commutes with expectation, the gradient term on the right side is unbiased:
$\mathbb{E}_{\tau,\epsilon}[-\nabla_{\theta}\ell^{w}_{\theta}(\tau,\epsilon)]=-\nabla_{\theta}\mathcal{L}_{\theta}^{w}{}(a_{t})=\nabla_{\theta}\text{ELBO}_{\theta}(a_{t}).$ (20)
In other words, gradient estimates are directionally unbiased even with worst-case overestimation of ratios. Our experiments are consistent with this result: while additional samples help, we observe empirically in Section 4.2 that FPO can be trained to outperform Gaussian PPO even with $N_{\text{mc}}=1$.
@@ 5 Discussion and Limitations
Acknowledgments
We thank Qiyang (Colin) Li, Oleg Rybkin, Lily Goli and Michael Psenka for helpful discussions and feedback on the manuscript. We thank Arthur Allshire, Tero Karras, Miika Aittala, Kevin Zakka and Seohong Park for insightful input and feedback on implementation details and the broader context of this work. This project was funded in part by NSF:CNS-2235013, IARPA DOI/IBC No. 140D0423C0035, and Bakar fellows. CK and BY are supported by NSF fellowship. SG is supported by the NVIDIA Graduate Fellowship